OpenStudy (anonymous):

A health club charges non-members $5 per day to swim and $9 per day for an exercise class. Members pay a yearly fee of $300 plus $4 per day to attend an exercise class and no swim fee. If Robert swims and takes an exercise class every time he goes to the gym, which equation shows the number of days he must use the gym to make the membership worthwhile?

OpenStudy (anonymous):

are there any answer choices?

OpenStudy (anonymous):

5d+9d/300=4d is the first one

OpenStudy (anonymous):

5d+9d/4d=300 is 2nd

OpenStudy (anonymous):

35d+9d = 300 - 4d and 5d+9d = 300 + 4d

OpenStudy (owllover101):

where is you smart score

OpenStudy (anonymous):

this may not help but I think its either the 1st or 2nd one